# Bayesian Hierarchical Probabilistic Forecasting of Intraday Electricity Prices

### v2

@@ Abstract
We present a first study of Bayesian forecasting of electricity prices traded on the German continuous intraday market which fully incorporates parameter uncertainty. A particularly large set of endogenous and exogenous covariables is used, handled through feature selection with Orthogonal Matching Pursuit (OMP) and regularising priors. Our target variable is the IDFull price index, forecasts are given in terms of posterior predictive distributions. For validation we use the exceedingly volatile electricity prices of 2022, which have hardly been the subject of forecasting studies before. As a benchmark model, we use all available intraday transactions at the time of forecast creation to compute a current value for the IDFull. According to the weak-form efficiency hypothesis, it would not be possible to significantly improve this benchmark built from last price information. We do, however, observe statistically significant improvement in terms of both point measures and probability scores. Finally, we challenge the declared gold standard of using LASSO for feature selection in electricity price forecasting by presenting strong statistical evidence that OMP leads to better forecasting performance.
@@ 1 Introduction and Motivation
In April 2024 the monthly share of renewable electricity generation in the European Union reached 52.3% (energy-charts.info), a new high and for the first time surpassing the milestone of having a majority of electrical energy produced by renewables in the EU. This milestone has already been reached March 2019 in Germany’s energy generation, now approaching a yearly share of 60% renewables. But the penetration of Renewable Energy Sources (RES) necessary for achieving the vision of sustainable energy production comes with tough challenges, as, opposed to other sources of energy, wind and solar energy production cannot be planned in advance, and mechanisms need to be in place to deal with the intermittency of RES in order to keep our energy systems stable.
@@
Figure 1: The unprecedented increase of volatility in the (German) market is illustrated. The volume-weighted average price (IDFull) for all hourly products for 2021 and 2022 are taken from EPEX Germany and shown as a thin line. The thick line is a weekly rolling average.
@@ 2 Electricity markets and market data
Figure 2: Illustration of the transformation proposed in Kulakov and Ziel (2019) to obtain the slope of the merit-order curve as a measure for elasticity $\eta$ in the DA market. The left side shows aggregated supply (SEP) and demand (DEM) curves as obtained from EPEX for the delivery hour 7:00 - 8:00 am on 4 November 2022, with the DA price $P_{\mathrm{da}}$ published in the European trading zone as a horizontal line. The right side depicts these curves after the transformation, in which the elasticity from the demand side is transferred to the supply side. Exemplary slopes are indicated by straight lines for two volumes marked by closed circles. A finite difference of 500 MWh has been used to determine the slopes.
@@ 2.1 Day-ahead market and elasticity
A subtle point in determining the DA market clearance price and volume is the different NEMOs and countries involved. Since 2014, the Single Day-Ahead Coupling (SDAC) creates a European trading zone for the DA market (nemo-committee.eu) and unifies the determination of the market clearance price through an algorithm called Euphemia (nemo-committee.eu). As a result, all participating countries and platforms, independent from the respective NEMO responsible, use the same SDAC price. The data published by EPEX since 2014 therefore includes only the unique SDAC price determined by Euphemia, but, on the other hand, is only authorised to publish the EPEX DA volume determined from EPEX auction curves in the respective country. From 14 October 2021, however, EPEX publishes the so-called All-Certified Exchanges’ aggregated supply and demand curves determined from the SDAC auction (EPEX Spot SE, 2023). The market clearance price and volume determined from the intersection of these curves will be the SDAC price and volume, while EPEX still publishes only their country-specific DA volume.
@@ 2.2 Continuous intraday market and transaction data
These market details are important in order to work with trading data from these markets. The full order book is rarely analysed in the literature due to the vast amount of data and a considerable increase of noise from automated trading, examples include Shinde et al. (2021); Scholz et al. (2021). In the present work, like in most of the CID literature, we have access to the executed transactions of the German side of all EPEX operated trades. These transactions include all trades within EPEX Germany (sell and buy side), and all SIDC trades where the sell or buy side is using the EPEX platform in Germany. The most important data fields available per transaction for the purpose of this study is the volume traded in MWh (Volume) with the matched price in EUR/MWh (Price), trade identification number (TradeId), the trade execution time (ExecutionTime), the begin of the delivery period (DeliveryStart), the end of the delivery period (DeliveryEnd), a flag if the transaction takes the buy or sell perspective (Side), and a flag indicating whether the sell and buy sides are identical (SelfTrade).
@@ 2.3 Reproduction of published EPEX CID price indices and statistics
As a general rule, the following conditions must be satisfied for a transaction to contribute to EPEX price indices and statistics (EPEX Spot SE, 2023).
@@
Criteria 1)-3) are met automatically by using the executed transaction data obtained from EPEX Spot. To ensure 4), we select transactions with SelfTrade=N (“No”) and SelfTrade=U (“Unknown”). The flag U can occur if one side of the transaction is on a trading platform operated by a different NEMO than EPEX (EPEX Spot SE, 2023). Naturally, these transactions can only be SIDC trades [^3] Note that not all SIDC trades need to be cross-border or cross-NEMO transactions, but, conversely, all cross-border or cross-NEMO transactions are the result of SIDC trades.. Condition 5) is directly ensured by comparing (DeliveryStart) and (DeliveryEnd) with the desired product [^4] Alternatively, one could use the Product information of transactions and select the products Intraday_Hour_Power and XBID_Hour_Power. However, occasionally, user-defined blocks of arbitrary delivery periods are traded which are listed as the closest product available, e.g. a 3 h delivery period would still be a Intraday_Hour_Power product.. Once the transactions are filtered to meet these conditions, duplicate entries have to be removed. The duplicates occur as EPEX includes all transactions executed on their platforms in the data, which includes the equivalent BUY and SELL sides if both sides are traded on the EPEX platform. To filter out these duplicates, we use the TradeId of each transaction, and only keep the transactions with the first occurring TradeId.
@@ 2.4 Live intraday values
Figure 3: Live CID price and volume indices and statistics calculated from transaction data for the exemplary delivery hour 7:00 - 8:00 am CET on 4 November 2022 as a function of forecast creation time $\tau$. The horizontal dashed lines indicate published values by EPEX. The dashed vertical line marks the beginning of delivery day $d$, i.e. hours left from this line are on day $d-1$ and on the right on day $d$. Solid vertical lines indicate 3 h before delivery start and 1 h before delivery start where applicable, and delivery start itself.
@@ 3 The data model
Day-ahead market information I EPEX Spot
@@
Continuous intraday market information I EPEX Spot
@@ 3.1 Market variables
Owing to the pan-European DA price, we only include the DA price of Switzerland. For Germany and SDAC countries, we also make use of DA price statistics provided by EPEX that aggregate DA prices of various hours of the day, e.g. morning, night, rush hour, sun peak; details are published by EPEX.
@@ 4 The forecast model
On a standard laptop, $N=140^{\prime}000$ samples $w_{i}$ and $\sigma_{i}$ of the parameter vector $(w,\sigma)^{\mathrm{T}}$ were obtained in matters of seconds for each forecast. With these samples, we estimate the PPD by substituting the posterior expectation with an average,
$\hat{p}_{\mathrm{ppd}}(y\,|\,X^{\ast},\bar{X},\bar{y})=\frac{1}{N}\sum_{i=1}^{N}$ $\,p_{\mathrm{like}}(y\,|\,X^{\ast},w_{i},\sigma_{i})$
$\quad\cdot p_{\mathrm{post}}(w_{i},\sigma_{i}\,|\,\bar{X},\bar{y}).$ (10)
@@ 5 Forecast study and discussion
XScenario (a), 7 July 2022 XX Scenario (e), 5 October 2022
Figure 7: Exemplary full day forecasts for all delivery hours $h$ of day $d=\;$7 July 2022 on the left and all delivery hours $h$ of day $d=\;$5 October 2022 on the right. The forecast on the left is part of scenario (a), where the forecast creation time $\tau$ is fixed to 23:00 pm on day $d-1$, implying an increasing forecast lag of $h_{\mathrm{lag}}=h+1$ along the horizontal axis. The forecast on the right is part of scenario (e), where the lag between forecast and delivery begin is fixed to 1 hour, i.e. $h_{\mathrm{lag}}=1$, implying forecast creation times $\tau=h-1$. The vertical colour bars represent the topology of the respective predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$ by means of $\alpha(p_{\mathrm{cut}})$ for 100 values of $p_{\mathrm{cut}}$, cf. (11).
@@ 5.2 Evaluation of point estimates
Figure 9: The top left chart depicts the difference between $\hat{y}$ and the live IDFull in terms of their MAE for scenarios (a) - (d), summarizing Figure 8. The bottom shows the same for scenarios (e) using OMP and (f) using LASSO for feature selection. Results below the solid red line possess smaller MAEs than their live IDFull references. In the top right chart, the MAE values are averaged across hours.
@@ 5.3 Spread sign forecasts
For an overall comparison of the sign forecast accuracy for both the spread and the rest, we investigate the choice $p_{0}=0.5$ for the fixed lag scenarios (e) and (f) in Figure 11, where the accuracy as the ratio of correct forecasts over total number of forecasts is shown. The forecasts using OMP feature selection results in a considerably higher accuracy than the forecasts using LASSO feature selection, and also turns out to be better than the benchmark $\mathrm{sgn}(P_{\mathrm{idfull}})$. Later we show that these results are statistically significant.
Figure 11: The accuracy of the estimators (16) and (17) for the sign of the spread and the rest are shown, together with the IDFull reference for the spread sign, cf. Figure 10. Here, the scenarios (e) and (f) are used to investigate the difference in performance using OMP and LASSO as feature selection. The accuracy is determined as number of correctly forecasted days across all hours over the total number of forecasts.
@@ 5.5 Statistical significance of results
Figure 14: Comparison between the use of OMP and LASSO for feature selection for scenarios (e) and (f) averaged across all days and hours of the test period. On the left we show the Average Coverage Error (ACE), on the right we show the Continuously Ranked Probability Score (CRPS) defined in (18).
Table 4: The $p$-values obtained by applying the one-sided Diebold-Mariano (DM) test are listed for various scores and forecast series’. As scores we consider the MAEs shown in the top right chart of Figure 9, a Boolean variable reflecting true sign forecasts of the spread and the rest as used for Figure 11, and the CRPS shown in Figure 14 on the right. The forecast series’ tested are those of scenarios (e) and (f), that is using OMP and LASSO for feature selection, as well as the live IDFull benchmark. Strong statistical significance is indicated by green colours, no statistical significance by red colours.
@@ 6 Conclusions
We presented a forecast study of the IDFull electricity price index of the German continuous intraday market for the years 2021 and 2022. These years are characterised by unprecedented volatility and have hardly been studied before. We provided details about reproducing price indices and statistics published by EPEX Spot using transaction data of the continuous intraday market.
@@
A currently debated topic in the literature is the supposed weak-form efficiency of the continuous intraday market. According to this hypothesis, all information available is already contained in last prices due to informed traders, making it impossible to significantly beat last price information by means of forecasting models. In our study, we partly confirm the hypothesis in that our model clearly identifies current price information as the dominating regressor and closely follows its trend. But on the other hand, we find that the live IDFull built from current prices can still be improved in a statistically significant way. It should, however, be taken into consideration that the definition of a last price benchmark is not unique, and other last price information, or combinations therefore, might still deliver the best forecast possible. It should also be noted that the continuous intraday market is developing rapidly, making it difficult for traders to adjust to changes, which in turn opens up the potential of forecast models to beat last price benchmarks.
Our conclusion on this debate from this study therefore is that the weak-form efficiency can tentatively be confirmed in the sense of a solid characterisation of market properties and possible future developments, but comprehensive forecast models may still be able to beat last price benchmarks. Aside from the question of weak-form efficiency of markets, last price benchmarks will still be point estimates, and even a probabilistic forecast that does not improve a last price benchmark but yields reliable uncertainty distributions around the benchmark is a fruitful directions of research and would be of great value for traders.

### v3

@@ Abstract
We address the need for forecasting methodologies that handle large uncertainties in electricity prices for continuous intraday markets by incorporating parameter uncertainty and using a broad set of covariables. This study presents the first Bayesian forecasting of electricity prices traded on the German intraday market. Endogenous and exogenous covariables are handled via Orthogonal Matching Pursuit (OMP) and regularising priors. The target variable is the IDFull price index, with forecasts given as posterior predictive distributions. Validation uses the highly volatile 2022 electricity prices, which have seldom been studied. As a benchmark, we use all intraday transactions at the time of forecast to compute a live IDFull value. According to market efficiency, it should not be possible to improve on this last-price benchmark. However, we observe significant improvements in point measures and probability scores, including an average reduction of $5.9\,\%$ in absolute errors and an average increase of $1.7\,\%$ in accuracy when forecasting whether the IDFull exceeds the day-ahead price. Finally, we challenge the use of LASSO in electricity price forecasting, showing that OMP results in superior performance, specifically an average reduction of $22.7\,\%$ in absolute error and $20.2\,\%$ in the continuous ranked probability score.
@@ 1 Introduction and Motivation
In April 2024 the monthly share of renewable electricity generation in the European Union reached 52.8% (energy-charts.info), a new high and for the first time surpassing the milestone of having a majority of electrical energy produced by renewables in the EU. This milestone has already been reached March 2019 in Germany’s energy generation, now approaching a yearly share of 60% renewables. But the penetration of Renewable Energy Sources (RES) necessary for achieving the vision of sustainable energy production comes with tough challenges, as, opposed to other sources of energy, wind and solar energy production cannot be planned in advance, and mechanisms need to be in place to deal with the intermittency of RES in order to keep our energy systems stable.
@@
Figure 1: The unprecedented increase of volatility in the (German) intraday electricity market is illustrated by the volume-weighted average price (IDFull) for all hourly products for 2021 and 2022 shown as a thin line. The thick line is a weekly rolling average. The data was provided by EPEX Spot SE .
@@ 2 Electricity markets and market data
Figure 2: Illustration of the transformation proposed in Kulakov and Ziel (2019) to obtain the slope of the merit-order curve as a measure for elasticity $\eta$ in the DA market. The left side shows aggregated supply (SEP) and demand (DEM) curves as obtained from EPEX Spot SE for the delivery hour 7:00 - 8:00 am on 4 November 2022, with the DA price $P_{\mathrm{da}}$ published in the European trading zone as a horizontal line. The right side depicts these curves after the transformation, in which the elasticity from the demand side is transferred to the supply side. Exemplary slopes are indicated by straight lines for two volumes marked by closed circles. A finite difference of 500 MWh has been used to determine the slopes.
@@ 2.1 Day-ahead market and elasticity
A subtle point in determining the DA market clearance price and volume is the different NEMOs and countries involved. Since 2014, the Single Day-Ahead Coupling (SDAC) creates a European trading zone for the DA market (nemo-committee.eu) and unifies the determination of the market clearance price through an algorithm called Euphemia (nemo-committee.eu). As a result, all participating countries and platforms, independent from the respective NEMO responsible, use the same SDAC price. The data published by EPEX since 2014 therefore includes only the unique SDAC price determined by Euphemia, but, on the other hand, is only authorised to publish the EPEX DA volume determined from EPEX auction curves in the respective country. From 14 October 2021, however, EPEX publishes the so-called All-Certified Exchanges’ aggregated supply and demand curves determined from the SDAC auction (EPEX Spot, 2023). The market clearance price and volume determined from the intersection of these curves will be the SDAC price and volume, while EPEX still publishes only their country-specific DA volume.
@@ 2.2 Continuous intraday market and transaction data
These market details are important in order to work with trading data from these markets. The full order book is rarely analysed in the literature due to the vast amount of data and a considerable increase of noise from automated trading, examples include Shinde et al. (2021); Scholz et al. (2021). In the present work, like in most of the CID literature, EPEX Spot SE provided us with the executed transactions of the German side of all EPEX operated trades. These transactions include all trades within EPEX Germany (sell and buy side), and all SIDC trades where the sell or buy side is using the EPEX platform in Germany. The most important data fields available per transaction for the purpose of this study is the volume traded in MWh (Volume) with the matched price in EUR/MWh (Price), trade identification number (TradeId), the trade execution time (ExecutionTime), the begin of the delivery period (DeliveryStart), the end of the delivery period (DeliveryEnd), a flag if the transaction takes the buy or sell perspective (Side), and a flag indicating whether the sell and buy sides are identical (SelfTrade).
@@ 2.3 Reproduction of published EPEX CID price indices and statistics
As a general rule, the following conditions must be satisfied for a transaction to contribute to EPEX price indices and statistics (EPEX Spot, 2023).
@@
Criteria 1)-3) are met automatically by using the executed transaction data obtained from EPEX Spot SE . To ensure 4), we select transactions with SelfTrade=N (“No”) and SelfTrade=U (“Unknown”). The flag U can occur if one side of the transaction is on a trading platform operated by a different NEMO than EPEX (EPEX Spot, 2023). Naturally, these transactions can only be SIDC trades [^3] Note that not all SIDC trades need to be cross-border or cross-NEMO transactions, but, conversely, all cross-border or cross-NEMO transactions are the result of SIDC trades.. Condition 5) is directly ensured by comparing (DeliveryStart) and (DeliveryEnd) with the desired product [^4] Alternatively, one could use the Product information of transactions and select the products Intraday_Hour_Power and XBID_Hour_Power. However, occasionally, user-defined blocks of arbitrary delivery periods are traded which are listed as the closest product available, e.g. a 3 h delivery period would still be a Intraday_Hour_Power product.. Once the transactions are filtered to meet these conditions, duplicate entries have to be removed. The duplicates occur as EPEX includes all transactions executed on their platforms in the data, which includes the equivalent BUY and SELL sides if both sides are traded on the EPEX platform. To filter out these duplicates, we use the TradeId of each transaction, and only keep the transactions with the first occurring TradeId.
@@ 2.4 Live intraday values
Figure 3: Live CID price and volume indices and statistics calculated from transaction data provided by EPEX Spot SE for the exemplary delivery hour 7:00 - 8:00 am CET on 4 November 2022 as a function of forecast creation time $\tau$. The horizontal dashed lines indicate published values by EPEX. The dashed vertical line marks the beginning of delivery day $d$, i.e. hours left from this line are on day $d-1$ and on the right on day $d$. Solid vertical lines indicate 3 h before delivery start and 1 h before delivery start where applicable, and delivery start itself.
@@ 3 The data model
Day-ahead market information I EPEX Spot SE
@@
Continuous intraday market information I EPEX Spot SE
@@ 3.1 Market variables
Owing to the pan-European DA price, we only include the DA price of Switzerland. For Germany and SDAC countries, we also make use of DA price statistics provided by EPEX Spot SE that aggregate DA prices of various hours of the day, e.g. morning, night, rush hour, sun peak; details are published by EPEX.
@@ 4 The forecast model
On a standard laptop, $N=140^{\prime}000$ posterior samples $w_{i}$ and $\sigma_{i}$ of the parameter vector $(w,\sigma)^{\mathrm{T}}$ were thus obtained in matters of seconds for each forecast. With these samples, we estimate the PPD by substituting the posterior expectation with an average,
$\hat{p}_{\mathrm{ppd}}(y\,|\,X^{\ast},\bar{X},\bar{y})=\frac{1}{N}\sum_{i=1}^{N}$ $\,p_{\mathrm{like}}(y\,|\,X^{\ast},w_{i},\sigma_{i}).$ (10)
@@ 5 Forecast study and discussion
XScenario (a), 6 July 2022 XX Scenario (e), 5 October 2022
Figure 7: Exemplary full day forecasts for all delivery hours $h$ of day $d=\;$6 July 2022 on the left and all delivery hours $h$ of day $d=\;$5 October 2022 on the right. The forecast on the left is part of scenario (a), where the forecast creation time $\tau$ is fixed to 23:00 pm on day $d-1$, implying an increasing forecast lag of $h_{\mathrm{lag}}=h+1$ along the horizontal axis. The forecast on the right is part of scenario (e), where the lag between forecast and delivery begin is fixed to 1 hour, i.e. $h_{\mathrm{lag}}=1$, implying forecast creation times $\tau=h-1$. The vertical colour bars represent the topology of the respective predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$ by means of $\alpha(p_{\mathrm{cut}})$ for 100 values of $p_{\mathrm{cut}}$, cf. (11).
@@ 5.2 Evaluation of point estimates
Figure 9: The top left chart depicts the difference between $\hat{y}$ and the live IDFull in terms of their MAE for scenarios (a) - (d), summarizing Figure 8. The bottom shows the same for scenarios (e) using OMP and (f) using LASSO for feature selection. Results below the solid red line possess smaller MAEs than their live IDFull references. In the top right chart, the MAE values are averaged across hours, with OMP showing an average decrease of $5.9\,\%$ in absolute errors compared to the live benchmark and a $22.7\,\%$ reduction compared to LASSO, while LASSO leads to an $28.8\,\%$ increase of absolute errors compared to the live benchmark.
@@ 5.3 Spread sign forecasts
For an overall comparison of the sign forecast accuracy for both the spread and the rest, we investigate the choice $p_{0}=0.5$ for the fixed lag scenarios (e) and (f) in Figure 11, where the accuracy as the ratio of correct forecasts over total number of forecasts is shown. The forecasts using OMP feature selection results in a considerably higher accuracy than the forecasts using LASSO feature selection, and also turns out to be better than the live benchmark $\mathrm{sgn}(P_{\mathrm{idfull}})$. Later we show that these results are statistically significant.
Figure 11: The accuracy of the estimators (16) and (17) for the sign of the spread and the rest are shown, together with the IDFull reference for the spread sign, cf. Figure 10. Here, the scenarios (e) and (f) investigate the difference in performance using OMP and LASSO for feature selection. OMP demonstrates an average increase of $1.7\,\%$ in spread sign forecast accuracy compared to the live benchmark, and increases of $12.1\,\%$ and $10.0\,\%$ for spread and rest sign forecast accuracy compared to LASSO. LASSO leads to a decrease in accuracy by $-9.0\,\%$ compared to the live benchmark. The accuracy is determined as number of correctly forecasted days across all hours over the total number of forecasts.
@@ 5.5 Statistical significance of results
Figure 14: Comparison between the use of OMP and LASSO for feature selection for scenarios (e) and (f) averaged across all days and hours of the test period. On the left, we show the Average Coverage Error (ACE), with OMP showing an average $34.59\,\%$ decrease compared to LASSO. On the right, we show the Continuous Ranked Probability Score (CRPS) defined in (18), with OMP showing an average $20.21\,\%$ decrease compared to LASSO.
Table 4: The average improvements and corresponding $p$-values obtained by applying the one-sided Diebold-Mariano (DM) test are listed for various scores and forecast series’. As scores we consider the MAEs shown in the top right chart of Figure 9, a Boolean variable reflecting true sign forecasts of the spread and the rest as used for Figure 11, and the CRPS shown in Figure 14 on the right. The forecast series’ tested are those of scenarios (e) and (f), that is using OMP and LASSO for feature selection, as well as the live IDFull benchmark. Strong statistical significance and improved forecast performance are indicated by green colours, no statistical significance and declined forecast performance by red colours.
@@ 6 Conclusions
We presented a forecast study of the IDFull electricity price index of the German continuous intraday market for the years 2021 and 2022. These years are characterised by unprecedented volatility and have hardly been studied before. We provided details about reproducing price indices and statistics published by EPEX Spot using transaction data of the continuous intraday market provided by EPEX Spot SE .
@@
A currently debated topic in the literature is the supposed weak-form efficiency of the continuous intraday market. According to this hypothesis, all information available is already contained in last prices due to informed traders, making it impossible to significantly beat last price information by means of forecasting models. In our study, we partly confirm the hypothesis in that our model clearly identifies current price information as the dominating regressor and closely follows its trend. But on the other hand, we find that the live IDFull built from current prices can still be improved in a statistically significant way. It should, however, be taken into consideration that the definition of a last-price benchmark is not unique, and other last price information, or combinations therefore, might still deliver the best forecast possible. It should also be noted that the continuous intraday market is developing rapidly, making it difficult for traders to adjust to changes, which in turn opens up the potential of forecast models to beat last-price benchmarks.
Our conclusion on this debate from this study therefore is that the weak-form efficiency can tentatively be confirmed in the sense of a solid characterisation of market properties and possible future developments, but comprehensive forecast models may still be able to beat last-price benchmarks. Aside from the question of weak-form efficiency of markets, last-price benchmarks will still be point estimates, and even a probabilistic forecast that does not improve a last-price benchmark but yields reliable uncertainty distributions around the benchmark is a fruitful directions of research and would be of great value for traders.